OpenStudy (anonymous):

Will give medal and a fan!!! How many 1/8's make 3/2?

OpenStudy (anonymous):

Please help

OpenStudy (amistre64):

well ... \[n*\frac{1}{8}=\frac32\] solve for n

OpenStudy (anonymous):

11/5?

OpenStudy (amistre64):

how would you solve for n?

OpenStudy (anonymous):

I dont know thats why i need help

OpenStudy (amistre64):

do you agree that 8/8=1?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

then lets multiply both side by 8 \[n*\frac 88 = 8*\frac 32\] \[n= \frac {8*3}2\] and simplify

OpenStudy (anonymous):

16 1/2?

OpenStudy (amistre64):

you might need to brush up on your multiplication and division skills

OpenStudy (mathmale):

I agree with this approach. Another approach would be to simply DIVIDE 3/2 by the fraction 1/8. Remember, when we divide a quantity by a fraction, we invert the fraction and multiply:\[\frac{ \frac{ 3 }{ 2 } }{ \frac{ 1 }{ 8 }}=\frac{ 3 }{ 2 }*\frac{ 8 }{ 1 }=?\]

OpenStudy (anonymous):

Oops i thought that said 8 x 8

OpenStudy (amistre64):

:) even 64/2 is not 16 1/2

OpenStudy (anonymous):

is it 12?

OpenStudy (anonymous):

Like is the answer 12?

OpenStudy (amistre64):

'we invert and multiply' is one of those methods (memory devices) that look good but which i feel add to the burden of simply not working the math ... like FOIL. I tend to prefer to multiply by a useful form of 1 \[\large \cfrac{ \frac{ 3 }{ 2 } }{ \frac{ 1 }{ 8 }}\] \[\large \cfrac{ \frac{ 3 }{ 2 } }{ \frac{ 1 }{ 8 }}\cfrac{ \frac{ 8 }{ 1 } }{ \frac{ 8 }{ 1 }}\] \[\large \cfrac{ \frac{ 3 }{ 2 }\frac{ 8 }{ 1 } }{ \frac{ 8 }{ 8 }}\] \[\large \cfrac{ \frac{ 3 }{ 2 }\frac{ 8 }{ 1 } }{1}\] \[ \frac{ 3 }{ 2 }\frac{ 8 }{ 1 } \] not that flip and multiply is bad advice, just seems like a memory device to me

OpenStudy (amistre64):

12 is good to me yes; half of 24

OpenStudy (anonymous):

it is 12

OpenStudy (anonymous):

Thank you so much @amistre64 and @kelliegirl33 and @mathmale!!

OpenStudy (amistre64):

i spose i should get to work and put all these labels on all those envelopes .... gotta love a desk job

OpenStudy (amistre64):

good luck ;)

OpenStudy (anonymous):

12